A certain two digit number is 4 more than 10 times the tens digit. The original number is 12 less than twice the number formed by interchanging the digits. Find the original number.
Answer by Boreal(15235)

10x+y original number
10y+x is interchanged
10x+y+12=2(10y+x)
10x+y=10x+4
y=4
10x+4+12=20y+2x
10x+16=80+2x
8x=64
x=8
the number is 84, and that is 4 more than 10 * 8.
interchanging the number yields 48, and 84 is 2 *48-12

84.